OpenStudy (anonymous):

graph the line 5x-2y=8

OpenStudy (anonymous):

First lets put it in slope-intercept form

OpenStudy (triciaal):

to graph a line you need 2 points. the easiest to use are when x = 0 what is y and when y = 0 what is x so you have when x = 0 -2y = 8 so y = -4 the point (0, -4) when y = 0 5x = 8 so x = 8/5 the point (8/5, 0) plot the points and connect y = 5/2 x -4

OpenStudy (anonymous):

You can graph a line without points. Just plot your y-intercept, then do up and over, or down and over from that point with the slope and you'll have a line.